# Singular flat bands in three dimensions: Landau level spreading, quantum geometry, and Weyl reconstruction

### v1

@@ III.2 Effective Weyl semimetal model
The Landau level spreading of the 3D singular flat band can be quantitatively understood using a band model that incorporates only the orbital Zeeman splitting, while neglecting Landau quantization. Specifically, we consider a modified Hamiltonian
@@
[…]
Most notably, the band structure in Fig. 4(a) exhibits a characteristic feature of a Weyl semimetal [66; 67; 68; 69; 70; 71], where the dispersive band and the upper flat band touch at two points in 3D momentum space, $\bm{k}_{\pm}=\pm\hat{\bm{z}}/(\sqrt{2}l_{B})$, where $\hat{\bm{z}}$ is a unit vector along $z$. In general, the energy bands of Weyl semimetals comprise a pair of band-touching points with opposite helicities. In the presence of a magnetic field, a chiral zeroth Landau level emerges, connecting the valence and conduction bands as illustrated in Fig. 4(c). In the present system, the chiral Landau level corresponds to the 0th level in the $j=-1$ sector, shown as the red flat line in Fig. 3, and schematically illustrated in Fig. 4(b) as a special case of the configuration in Fig. 4(c). In this context, we conclude that the presence of the special flat Landau level in Fig. 3 is a consequence of the Weyl semimetal formed by the orbital Zeeman splitting.
@@ III.3 Quantum geometry and Landau level spreading
In two-dimensional singular flat-band systems such as the kagome lattice, the Landau level energy spread is known to be quantitatively linked to the underlying quantum geometry of the energy band [55; 4; 5]. This raises the intriguing question of whether a similar geometric framework can be extended to 3D flat band systems. In our 3D case, with a magnetic field applied along the $z$-direction, the system can be regarded as a collection of two-dimensional subsystems indexed by $k_{z}$. Unlike the 2D counterparts [55; 4; 48], however, the flat bands in this system remain doubly degenerate throughout the Brillouin zone, where the non-Abelian nature of the quantum geometric tensor [63; 64; 65] complicates the direct application of existing Abelian-based approaches.
@@
In the limit of $k_{z}\rightarrow\infty$, the Hamiltonian is dominated by the zeroth-order term $H_{\mathrm{eff+Z}}^{(0)}$, which depends solely on the orbital angular momentum operator $L_{z}$. As a result, the eigenstates of the Hamiltonian are simply the eigenstates of $L_{z}$, expressed as $|\psi^{(0)}_{m}\rangle=\vec{e}_{m}$, where $m=-1,0,1$ denotes the eigenvalue of $L_{z}$. In this regime, the eigenstates become independent of $k_{x}$ and $k_{y}$, indicating a flat geometry in the $k_{x}$$k_{y}$ plane. Correspondingly, the director shown in Fig.1(b) aligns uniformly as $k_{z}\rightarrow\infty$. This trivial quantum metric reflects the vanishing Landau level spreading in the large-$k_{z}$ limit, as seen in Fig.3.
@@
$|\psi_{\pm 1}\rangle=\frac{1}{\sqrt{N_{\pm 1}}}\left(\exp(\mp i\phi_{k})\vec{e}_{\pm 1}\mp\frac{1}{\sqrt{2}}\frac{k_{\parallel}k_{z}l_{B}^{2}}{k_{z}^{2}l_{B}^{2}\mp 1/2}\vec{e}_{0}\right),$ (43)
where $N_{\pm 1}=1+k_{\parallel}^{2}k_{z}^{2}l_{B}^{4}/[2(k_{z}^{2}l_{B}^{2}\mp 1/2)^{2}]$. From these expressions, the quantum metric component along the azimuthal direction can be computed as
@@
$\simeq$ $\sqrt{2}\left(k_{\parallel}/k_{z}\right)^{2}$ (44)
to leading order in $k_{\parallel}/k_{z}$. In the large $k_{z}$ regime, the Landau level spreading is given by the second term of Eq. (39), and it is proprtional to $j/k_{z}^{2}$ with the total angular momentum index $j$. The metric component of Eq. (III.3) is proportional to this band width, noting that $j\sim n\sim k_{\parallel}^{2}$ for large $n\gg 1$. This establishes a clear link between the quantum metric properties of the Zeeman-split flat band and calculated Landau level spreading. At $k_{z}\lesssim k_{\parallel}$, the degenerate flat bands are strongly hybridized.
@@ IV Shubnikov-de Haas Oscillations in bent flat bands
where we have shifted the energy zero point by $-12t^{\prime}$. Accordingly, the continuum zero-field Hamiltonian [Eq. (6)] and the Landau-level Hamiltonian [Eq. (III.1)] acquire correction terms,
@@
$H^{\prime}_{B}=-\frac{8t^{\prime}}{t}\hbar\omega_{B}\left(a^{\dagger}a+\frac{1}{2}+k_{z}^{2}l_{B}^{2}\right)$ (47)
@@
Figure 5(a) shows the Landau levels in the flat band region, calculated from $H_{\mathrm{B}}+H_{\mathrm{B}}^{\prime}$ with $8t^{\prime}=0.005t$. For $t^{\prime}>0$, the Landau levels bend downward with increasing $k_{z}$. In Fig. 5(b), we show the corresponding energy band of the effective model with the orbital Zeeman term, $H_{\mathrm{eff+Z}}=H_{\mathrm{eff}}+H_{\mathrm{eff}}^{\prime}+H_{Z}$, where the Weyl points are indicated by red spheres. We observe good agreement between the Landau level structure and the Weyl semimetal band even in the presence of the additional band dispersion. Here we note that the momentum $k_{z}$ (horizontal axis) and the energy $E$ (vertical axis) are scaled by $l_{B}$ and $\hbar\omega_{B}$, respectively, so that the plots in Figs. 5(a) and (b) are independent of the magnetic field amplitude under this scaling.
Figure 5: Electronic structure of slightly bended flat band under the magnetic field. (a) Landau level structure. (b) Dispersion relation of $H_{\mathrm{eff}}+H_{\mathrm{eff}}^{\prime}+H_{Z}$, with Weyl points indicated by red spheres. The inset shows a three-dimensional plot of (b) as a function of $k_{z}$ and $k_{\parallel}$. The parameter $0.0058t^{\prime}/t=$ 0.005 is used.
@@
As the magnetic field increases, the Zeeman splitting becomes more pronounced, and the Fermi surface evolves from two degenerate spheres [Fig. 7(i)] to split lobes along the $k_{z}$ direction [Fig. 7(ii)]. With further increase in $B$, the Fermi surfaces become fully separated and elongated along the $k_{z}$ axis, leading to a reduced cross-sectional area in the $k_{x}$$k_{y}$ plane [Fig. 7(vi)]. This evolution directly influences the SdH peak spacing. In Fig. 6(c), we overlay the plot of $S_{\mathrm{F}}^{\perp}$ as a function of inverse magnetic field, where we observe qualitative agreement with the spacing, evaluated from $(B_{n+1}^{-1}-B_{n}^{-1})$. In addition to the overall nonlinear trends associated with Fermi surface evolution, two shoulder-like features appear at the vertical lines in Fig. 6(c). These correspond to the magnetic field $B$ where the Fermi energy intersects the minimum of the first Landau level [$\nu=+1$, yellow curve in Fig. 5(a)] and the top of the second Landau levels ($\nu=-1$). These results indicate that the characteristic Landau level structure of the singular flat bands can be identified from the non-uniform spacing of the SdH oscillations.
@@ V Extention to High Angular Momentum flat bands
where $\bm{L}^{(\ell)}=(L_{x}^{(\ell)},L_{y}^{(\ell)},L_{z}^{(\ell)})$ are the set of angular momentum matrices for a given $\ell$, satisfying $(\bm{L}^{(\ell)})^{2}=\ell(\ell+1)$. The explicit matrix expression is provided in Appendix E. The eigenenergies are given by $E_{m}=(td^{2}/2)(\ell^{2}-m^{2})k^{2}$, where $m=-\ell,-\ell+1,\cdots,\ell$ is the magnetic quantum number. Notably, the states with $m=\pm\ell$ become exactly flat with zero energy, $E_{\pm\ell}=0$, yielding a doubly degenerate flat band. The remaining $12\ell-$ 1 bands are dispersive and parabolic, touching the flat band from above. The case of $\ell=1$ corresponds to the continuum model for the pyrochlore lattice introduced in Eq. (19). In addition, $\ell=3/2$ can be viewed as the Luttinger-Kohn model [72], $H_{\mathrm{LK}}=A\bm{k}^{2}+B(\bm{k}\cdot\bm{L}^{(3/2)})^{2}$, with a special parameter choice $A=-(3/2)^{2}B$.
@@
which is a matrix of dimension $12\ell+$ 1. Here $\tilde{H}_{B}(j,k_{z})$ is structured as
@@
with $L_{z}=\mathrm{diag}(\ell,\ell-1,\cdots,-\ell)$ is naturally incorporated and leads to a splitting of levels according to the angular momentum projection $m$. For sectors with $-\ell\leq j<\ell$, the projected Hamiltonian is obtained as the lower-right $(\ell+j+1)\times(\ell+j+1)$ submatrix of $H_{B}(j,k_{z})$, following the same construction as described in Sec. III.
Figure 8(a) shows the energy spectrum for the Landau level Hamiltonian $H_{B}$ for $\ell=3/2,2$ and 5/2. The Landau levels originating from the flat band are spread over the range $-\Delta_{Z}<E<\Delta_{Z}$, where
@@
indicating that the spread becomes wider with increasing $\ell$. The $j=-\ell$ level, shown as a red line, remains perfectly flat as a function of $k_{z}$, and also isolated from the main cluster the Landau levels. Notably, for arbitrary $\ell$, the Landau level spreading in the flat band is completely suppressed in the reduced Hamiltonian $\tilde{H}_{B}=H_{B}-H_{\mathrm{Z}}$, as already observed in the case of $\ell=1$ [Fig. 3(b)]. This can be shown by noting that the reduced Hamiltonian $\tilde{H}_{B}$ [Eq. (52)] has rank $12\ell-$ 1, which is two less than its full dimension $12\ell+$ 1, implying the existence of two zero modes. We thus conclude that, regardless of the angular momentum $\ell$, the Landau level spreading originates solely from the Zeeman term $H_{\mathrm{Z}}$.
@@
The Landau level spectrum reflects the band structure of the effective model with the Zeeman term, $H_{\mathrm{eff}+Z}=H_{\mathrm{eff}}+H_{\mathrm{Z}}$, which is shown in Fig. 8(b). Importantly, there are $12\ell-$ 1 Weyl points on the $k_{z}$ axis, which correspond to the band inversion of the states with different angular momentum $m$, as schematically shown in Fig. 9(a). As $k_{z}$ increases, the $m=-\ell$ band (the flat band) crosses the $m=-\ell+1$, $-\ell+2$, $\cdots$, $\ell-1$ bands. Each band inversion between $m$ and $m^{\prime}$ yields a Weyl point with helicity $\chi=m^{\prime}-m$, so the system hosts Weyl points with $\chi=1$, $2$, $3$, $\cdots$ $12\ell-$ 1.

### v2

@@ III.2 Effective Weyl semimetal model
In conventional electronic systems, the Landau level spectrum can be approximately understood via semiclassical Bohr–Sommerfeld quantization, in which Landau levels emerge from the discretization of constant-energy contours of the original energy band. In singular flat-band systems, however, this semiclassical framework does not straightforwardly account for the observed Landau level spreading, as a dispersionless flat band leads to degenerate Landau levels at a single energy within this scheme. In this section, we show that a band model incorporating only the orbital Zeeman effect provides an effective underlying continuous band dispersion that accounts for the spread of the Landau levels.
Specifically, we consider a modified Hamiltonian
@@
in which the orbital Zeeman term $H_{\mathrm{Z}}$ in Eq. (35) is added to the continuum Hamiltonian $H_{\mathrm{eff}}$ in Eq. (19), without replacing the momentum $\bm{k}$ with $\bm{\pi}/\hbar$. Figure 4(a) shows the band dispersion of $H_{\mathrm{eff+Z}}$ as a function of $k_{z}$ and $k_{\parallel}$, where $k_{\parallel}$ denotes the in-plane momentum along an arbitrary direction in the $k_{x}k_{y}$ plane. The resulting band structure reproduces the overall spectral features of the Landau levels as a function of $k_{z}$ shown in Fig. 3, except for the discrete level structure arising from the quantization of the $k_{x}k_{y}$ plane. In particular, the 3D flat band splits into $\pm\hbar\omega_{B}/4$ in the regime $k_{z}l_{B}\gg 1$, and the upper flat-band branch becomes dispersionless at $k_{z}=0$.
Most notably, the band structure in Fig. 4(a) exhibits a characteristic feature of a Weyl semimetal [66, 67, 68, 69, 70, 71], where the dispersive band and the upper flat band touch at two points in 3D momentum space, $\bm{k}_{\pm}=\pm\hat{\bm{z}}/(\sqrt{2}l_{B})$, where $\hat{\bm{z}}$ is a unit vector along $z$. In general, the energy bands of Weyl semimetals comprise a pair of band-touching points with opposite helicities. In the presence of a magnetic field, a chiral zeroth Landau level emerges, connecting the valence and conduction bands as illustrated in Fig. 4(c). In the present system, the chiral Landau level corresponds to the 0th level in the j=1j=-1 sector, shown as the red flat line in Fig. 3, and schematically illustrated in Fig. 4(b) as a special case of the configuration in Fig. 4(c).
To better illustrate this correspondence, we introduce a modified model in which the Zeeman term is artificially enhanced. Specifically, we consider the continuum and quantized Landau level Hamiltonians defined as
$H_{\mathrm{eff}+\eta\mathrm{Z}}=H_{\mathrm{eff}}(\bm{k})+\eta H_{\mathrm{Z}},$ (41)
$H_{B,\eta}(j,k_{z})=\tilde{H}_{B}+\eta H_{\mathrm{Z}},$ (42)
where $\eta$ is an artificial enhancement factor applied to the orbital Zeeman term. The calculated band structures for $\eta=50$, obtained from Eqs. (41) and (42), are shown in Figs. 5(a) and 5(b), respectively. The enhancement factor $\eta$ effectively rescales the bandwidth of the originally flat-band cluster, while the Landau level spacing in the quantized model [Fig. 5(b)] remains essentially unchanged. In Fig. 5(b), we clearly observe that the chiral Landau level (red curve) penetrates the two Weyl points of the continuum spectrum shown in Fig. 5(a). As $\eta$ is reduced toward $\eta=1$, the spectrum continuously evolves back to that of Fig. 4(b), while retaining its key spectral features. This analysis supports our conclusion that the emergence of the special flat Landau level in Fig. 3 is a direct consequence of the Weyl semimetal structure induced by the orbital Zeeman effect.
Figure 5: Comparison between the energy spectrum of the effective Weyl model and original Landau level structure, calculated with an enhanced Zeeman term. The enhancement factor is taken to be $\eta=50$.
@@ III.3 Quantum geometry and Landau level spreading
In two-dimensional singular flat-band systems such as the kagome lattice, the Landau level energy spread is known to be quantitatively linked to the underlying quantum geometry of the energy band [55, 4, 5]. This raises the intriguing question of whether a similar geometric framework can be extended to 3D flat band systems. In our 3D case, with a magnetic field applied along the $z$-direction, the system can be regarded as a collection of two-dimensional subsystems indexed by $k_{z}$. Unlike the 2D counterparts [55, 4, 48], however, the flat bands in this system remain doubly degenerate throughout the Brillouin zone, where the non-Abelian nature of the quantum geometric tensor [63, 64, 65] complicates the direct application of existing Abelian-based approaches.
@@
In the limit of $k_{z}\rightarrow\infty$, the Hamiltonian is dominated by the zeroth-order term $H_{\mathrm{eff+Z}}^{(0)}$, which depends solely on the orbital angular momentum operator $L_{z}$. As a result, the eigenstates of the Hamiltonian are simply the eigenstates of $L_{z}$, expressed as $|\psi^{(0)}_{m}\rangle=\vec{e}_{m}$, where m=1,0,1m=-1,0,1 denotes the eigenvalue of $L_{z}$. In this regime, the eigenstates become independent of $k_{x}$ and $k_{y}$, indicating a flat geometry in the $k_{x}$$k_{y}$ plane. Correspondingly, the director shown in Fig.1(b) aligns uniformly as $k_{z}\rightarrow\infty$. This trivial quantum metric reflects the vanishing Landau level spreading in the large-$k_{z}$ limit, as seen in Fig.3.
@@
$|\psi_{\pm 1}\rangle=\frac{1}{\sqrt{N_{\pm 1}}}\left(\exp(\mp i\phi_{k})\vec{e}_{\pm 1}\mp\frac{1}{\sqrt{2}}\frac{k_{\parallel}k_{z}l_{B}^{2}}{k_{z}^{2}l_{B}^{2}\pm 1/2}\vec{e}_{0}\right),$ (45)
where $N_{\pm 1}=1+k_{\parallel}^{2}k_{z}^{2}l_{B}^{4}/[2(k_{z}^{2}l_{B}^{2}\pm 1/2)^{2}]$. From these expressions, the quantum metric component along the azimuthal direction can be computed as
@@
$\simeq$ $\frac{1}{\sqrt{2}}\left(k_{\parallel}/k_{z}\right)^{2}$ (46)
to leading order in $k_{\parallel}/k_{z}$. Notably, this component is directly proportional to the maximum Hilbert Schmidt quantum distance between states on a circle at fixed $k_{z}$ and $k_{\parallel}$, which is given by
$d_{\mathrm{max},\pm}$ $=\max_{\phi_{k},\phi_{k^{\prime}}}(1-|\langle\psi_{\pm 1}(\phi_{k^{\prime}})|\psi_{\pm}(\phi_{k})\rangle|)$
$\simeq(k_{\parallel}/k_{z})^{2}=\sqrt{2}{g_{\phi_{k}\phi_{k},\pm}}$ (47)
In the large $k_{z}$ regime, the Landau level spreading is given by the second term of Eq. (39), and it is proprtional to $j/k_{z}^{2}$ with the total angular momentum index $j$. The metric component of Eq. (III.3), as well as the maximum quantum distance in Eq. (III.3), is proportional to this band width, noting that $j\sim n\sim k_{\parallel}^{2}$ for large $n\gg 1$. This establishes a clear link between the quantum metric properties of the Zeeman-split flat band and calculated Landau level spreading. At $k_{z}\lesssim k_{\parallel}$, the degenerate flat bands are strongly hybridized.
@@ IV Shubnikov-de Haas Oscillations in bent flat bands
where we have shifted the energy zero point by $12t^{\prime}$. Accordingly, the continuum zero-field Hamiltonian [Eq. (6)] and the Landau-level Hamiltonian [Eq. (III.1)] acquire correction terms,
@@
$H^{\prime}_{B}=-\frac{4t^{\prime}}{t}\hbar\omega_{B}\left(a^{\dagger}a+\frac{1}{2}+k_{z}^{2}l_{B}^{2}\right)$ (50)
@@
Figure 6(a) shows the Landau levels in the flat band region, calculated from $H_{\mathrm{B}}+H_{\mathrm{B}}^{\prime}$ with $4t^{\prime}=0.0025t$. For $t^{\prime}>0$, the Landau levels bend downward with increasing $k_{z}$. In Fig. 6(b), we show the corresponding energy band of the effective model with the orbital Zeeman term, $H_{\mathrm{eff+Z}}=H_{\mathrm{eff}}+H_{\mathrm{eff}}^{\prime}+H_{Z}$, where the Weyl points are indicated by red spheres. We observe good agreement between the Landau level structure and the Weyl semimetal band even in the presence of the additional band dispersion. Here we note that the momentum $k_{z}$ (horizontal axis) and the energy $E$ (vertical axis) are scaled by $l_{B}$ and $\hbar\omega_{B}$, respectively, so that the plots in Figs. 6(a) and (b) are independent of the magnetic field amplitude under this scaling.
Figure 6: Electronic structure of slightly bended flat band under the magnetic field. (a) Landau level structure. (b) Dispersion relation of $H_{\mathrm{eff}}+H_{\mathrm{eff}}^{\prime}+H_{Z}$, with Weyl points indicated by red spheres. The inset shows a three-dimensional plot of (b) as a function of $k_{z}$ and $k_{\parallel}$. The parameter $4t^{\prime}/t=0.0025$ is used.
@@
As the magnetic field increases, the Zeeman splitting becomes more pronounced, and the Fermi surface evolves from two degenerate spheres [Fig. 8(i)] to split lobes along the $k_{z}$ direction [Fig. 8(ii)]. With further increase in $B$, the Fermi surfaces become fully separated and elongated along the $k_{z}$ axis, leading to a reduced cross-sectional area in the $k_{x}$$k_{y}$ plane [Fig. 8(vi)]. This evolution directly influences the SdH peak spacing. In Fig. 7(c), we overlay the plot of $S_{\mathrm{F}}^{\perp}$ as a function of inverse magnetic field, where we observe qualitative agreement with the spacing, evaluated from $(B_{n+1}^{-1}-B_{n}^{-1})$. In addition to the overall nonlinear trends associated with Fermi surface evolution, two shoulder-like features appear at the vertical lines in Fig. 7(c). These correspond to the magnetic field $B$ where the Fermi energy intersects the minimum of the first Landau level $\nu=+1$, yellow curve in Fig. 6(a)] and the top of the second Landau levels ($\nu=-1$). These results indicate that the characteristic Landau level structure of the singular flat bands can be identified from the non-uniform spacing of the SdH oscillations.
@@ V Extention to High Angular Momentum flat bands
where $\bm{L}^{(\ell)}=(L_{x}^{(\ell)},L_{y}^{(\ell)},L_{z}^{(\ell)})$ are the set of angular momentum matrices for a given $\ell$, satisfying $(\bm{L}^{(\ell)})^{2}=\ell(\ell+1)$. The explicit matrix expression is provided in Appendix E. The eigenenergies are given by $E_{m}=(td^{2}/2)(\ell^{2}-m^{2})k^{2}$, where $m=-\ell,-\ell+1,\cdots,\ell$ is the magnetic quantum number. Notably, the states with $m=\pm\ell$ become exactly flat with zero energy, $E_{\pm\ell}=0$, yielding a doubly degenerate flat band. The remaining $2\ell-1$ bands are dispersive and parabolic, touching the flat band from above. The case of $\ell=1$ corresponds to the continuum model for the pyrochlore lattice introduced in Eq. (19). In addition, $\ell=3/2$ can be viewed as the Luttinger-Kohn model [72], $H_{\mathrm{LK}}=A\bm{k}^{2}+B(\bm{k}\cdot\bm{L}^{(3/2)})^{2}$, with a special parameter choice $A=-(3/2)^{2}B$.
@@
which is a matrix of dimension $2\ell+1$. Here $\tilde{H}_{B}(j,k_{z})$ is structured as
@@
with $L_{z}=\mathrm{diag}(\ell,\ell-1,\cdots,-\ell)$ is naturally incorporated and leads to a splitting of levels according to the angular momentum projection $m$. For sectors with $\ell\leq j<\ell$, the projected Hamiltonian is obtained as the lower-right $(\ell+j+1)\times(\ell+j+1)$ submatrix of $H_{B}(j,k_{z})$, following the same construction as described in Sec. III.
Figure 9(a) shows the energy spectrum for the Landau level Hamiltonian $H_{B}$ for $\ell=3/2,2$ and 5/2. The Landau levels originating from the flat band are spread over the range $\Delta_{Z}<E<\Delta_{Z}$, where
@@
indicating that the spread becomes wider with increasing $\ell$. The $j=$-$\ell$ level, shown as a red line, remains perfectly flat as a function of $k_{z}$, and also isolated from the main cluster the Landau levels. Notably, for arbitrary $\ell$, the Landau level spreading in the flat band is completely suppressed in the reduced Hamiltonian $\tilde{H}_{B}=H_{B}-H_{\mathrm{Z}}$, as already observed in the case of $\ell=1$ [Fig. 3(b)]. This can be shown by noting that the reduced Hamiltonian $\tilde{H}_{B}$ [Eq. (55)] has rank $2\ell-1$, which is two less than its full dimension $2\ell+1$, implying the existence of two zero modes. We thus conclude that, regardless of the angular momentum $\ell$, the Landau level spreading originates solely from the Zeeman term $H_{\mathrm{Z}}$.
@@
The Landau level spectrum reflects the band structure of the effective model with the Zeeman term, $H_{\mathrm{eff}+Z}=H_{\mathrm{eff}}+H_{\mathrm{Z}}$, which is shown in Fig. 9(b). Importantly, there are $2\ell-1$ Weyl points on the $k_{z}$ axis, which correspond to the band inversion of the states with different angular momentum $m$, as schematically shown in Fig. 10(a). As $k_{z}$ increases, the $m=$-$\ell$ band (the flat band) crosses the $m=-\ell+1$, $\ell+2$, $\cdots$, $\ell-1$ bands. Near each crossing point, the low energy physics is described by a two-band effective model involving the $m=$-$\ell$ and $m^{\prime}$ states. The corresponding Weyl Hamiltonian takes the form
$H_{\mathrm{W}}=\Delta_{Z}+\left(\begin{array}[]{cc}0&Ak_{+}^{\ell+m^{\prime}}\\ Ak_{-}^{\ell+m^{\prime}}&v_{m^{\prime}}(k_{z}-k_{\mathrm{W}})\end{array}\right)$ (61)
where $k_{+}=k_{x}+ik_{y}$, $A$ is a real parameter and $v_{m^{\prime}}$ is the group velocity of of the $m^{\prime}$ state along the $z$ direction at the Weyl point at $\bm{k}=k_{\mathrm{W}}\hat{\bm{z}}$. In this system, velocity satisfies $v_{m^{\prime}}k_{\mathrm{W}}>0$. Consequently, each band inversion between $m=$-$\ell$ and $m^{\prime}$ yields a Weyl point with helicity $\chi=\pm(m-m^{\prime})$ depending on whether $k_{\mathrm{W}}\gtrless 0$. The system therefore hosts Weyl points with $\chi=\pm 1$, $\pm 2$, $\pm 3$, $\cdots$ $\pm(2\ell-1)$, where each helicity corresponds to the Chern number defined on a sphere enclosing the respective Weyl point.